Find all solutions, real and complex, of the equation. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.)
x4 − 625 = 0

Answer by lwsshak3(11628)

Find all solutions, real and complex, of the equation. (Enter your answers as a comma-separated list. If there is no solution, enter NO SOLUTION.)
x4 − 625 = 0
**
x4−625 = (x^2+25)(x^2-25)= (x^2+25)(x+5)(x-5)=0
..
x^2+25=0
x^2=-25
x=±√-25=±5i
..
x+5=0
x=-5
..
x-5=0
x=5
..
solutions: -5i, 5i, -5, 5,